

********************************************************************
UARS WEEKLY STATUS REPORT - 05 December 1997
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
***********
PEM
***********
PEM/AXIS was active from November 26 through December 4. PEM zenith
spectrometers were activated on December 4. PEM produced energy deposition
profiles from electrons inverted from the AXIS X-ray measurements and
is currently generating profiles of the electron and proton precipitation
from spectrometer data.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
HALOE was power on December 3, 1997, a day earlier than scheduled to
provide correlative data to the shuttle instrument SOLSE. All instrument
parameters are nominal and the science data is being processed.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI is currently off. Next turn-on is scheduled for Dec. 8.
Next week Prof. Paul Hays and Dr. David Ortland will be attending the Fall
AGU meeting in San Francisco to present the following papers.
"HRDI: A Study of Dynamics"
"Analysis of temperatures and winds in the mesosphere and lower
thermosphere measured by HRDI"
***********
WINDII
***********
WINDII has been operating since November 24, and is operating
with no problems. We can't say "normally", as it is running
with the active temperature control system off, in order to
reduce the power requirements, and particularly the warmup
surge that would otherwise occur. The temperature seems rather
stable, even without active control.
Five WINDII team members are attending the AGU meeting, and
contributing to the UARS Aeronomy session.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 350 CD-ROM requests as of 05-December-97. One
new CD-ROM order was received and two were shipped this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
20 days PEM (3.32) Prod (1350R-635C=715L)
The CDHF continues to monitor and query the catalog for UKMO data gaps.
The catalog is now up to date except for UARS DAYS 2263 and 2264 and we have
requested that these days be sent.
The CDHF supported CLAES calibration tables and other file management issues
related to their upcoming delivery.
The CDHF will be implementing line tests to RACs with poor performance, in an
effort to log network throughput to various sites.
Pending software upgrades: IDL 5.0, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 November 1997 (Orbit 33864, GMT Day 325) through 28
November 1997 (Orbit 33967, GMT Day 332).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 328/0124 33894 ZEPS off
WINDII 328/0631 33897 On
WINDII 328/0816 33898 Off
WINDII 328/1600 33903 On
HRDI 329/2059 33921 Off
PEM 330/0139 33924 AXIS on
HRDI 330/1715 33934 Off, low
The beta angle reached a maximum angle of 78.1 degrees on 27 November 1997 (DOY
331). The beta angle is now decreasing toward a minimum angle of 0.0 degrees
on 17 December 1997 (DOY 351).
Relative Time Sequences (RTSs) 12 and 13 were modified and uplinked to the
spacecraft on 26 November 1997 (see Other Significant Events section) to
correct the compensation heater cycling scheme.
The WINDII turn-on procedure on 23 November 1997 was interrupted due to
unavailability of a reference microprocessor load (see Ground System section).
The problem was investigated, rectified, and WINDII turn-on was successfully
completed on 24 November 1997 (328/1600).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 27.2 end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 degrees C. Additional battery performance data for each battery on
22 November 1997 and 28 November 1997 are provided in Appendix D.
The clock error ranged from -9.8 to +10.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 02
% OUT-OF-LIMITS = 10
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 04
% HRDI 03
% PEM 02
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT 04
TRANSPONDER FREQUENCIES:
Transponder A: 97/301 - 2287.497086
Transponder B: 97/302 - 2287.499098
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 97/331 show the observatory data loss to be 20 hours, 41
minutes, 02.292 seconds (no change). This is a 0.0382 percent data loss which
equals a 99.9618 percent data capture for the mission.
Relative Time Sequence (RTS) 13 was updated to include heater disable commands
for compensation heaters 19, 20, and 23. RTS 12 was modified to for
compensation heater 18Us A side turn on instead of having both the A and B side
turn on as in the previous version.
The WIRELOAD procedure used to turn-on WINDII Rhung-upS when the reference
PWIHCNTL file could not be found on the system. Investigation revealed that
the FOT was not informed that new microloads were submitted the previous week.
WINDII operations was contacted and the situation resolved. New load files
were generated and placed on the AP Prime Pack. The WIRELOAD procedure was
updated to match the filenames of the new loads submitted by WINDII.
Modifications to the OBCSTF resident POCC database were initiated this week.
Several minor modifications are planned and a new database (version 65) will be
tested and released before the next yaw maneuver scheduled for 17 December 1997.
The FOT is still working out the details of participating in the IP return-link
telemetry test that the network is conducting.
Flight Dynamics Facility (FDF) continues the generation process for a new Star
Catalog to fix the invalid precession rate for star 48000001.
No testing of the new CMS software release (R7B1V2) occurred this week.
Investigation of the stored command verification problem - time to complete -
associated with Applications Processor (AP) software release 10.08 continued.
The cause has been identified, but rectifying the problem continues.
OTHER SIGNIFICANT EVENTS
During power-off of HRDI (due to low power) on 25 November 1997, the FOT
noticed that the HRDI compensation heater (#19) was not cycling properly.
Further investigation showed that the HALOE compensation heater (#20) was also
not cycling properly (both were being kept on during spacecraft solar array
night). These heaters were supposed to be turned off at end-of-day via PMON
execution of RTS-13, which contains operational and compensation heater ROFFS
commands. More investigation then showed that the latest version of RTS-13
did not have the compensation heater off commands for heaters 19, 20, and 23
(HRDI, HALOE, and WINDII). RTS-12 (holding the heater RONS commands) was also
discovered to contain an additional, unnecessary, heater command for
compensation heater #18 (SUSIM).
Apparently, the correct versions of RTS-12 and RTS-13 were not reloaded
following simultaneous operation of MLS and HALOE in late October. During this
time, the SSPP instruments were powered off, and heater control for these
instruments was performed via several different versions of RTS-12 and RTS-13.
However, when the SSPP instruments were powered back on, the original versions
of RTS-12 and RTS-13 were not reloaded. As there is no way to verify via
telemetry which RTS version is on-board, the discrepancy was not noted.
Unfortunately, the faulty RTSs resulted in slightly increased night heater
load, which was also not noted, because of the large number of other
operational changes taking place. These heater loads probably account for the
difficulty the FOT has experienced since the end of October in meeting the
planned instrument schedule.
The correct versions of RTS-12 and RTS-13 were reloaded immediately upon
discovery of the missing commands on Tuesday, 25 November 1997 (329/22:36 UTC,
Orbit 33922). Recovery of additional power was immediate, and the power
subsystem has been stable since then.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current status
of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 17 December 1997
at 311/132708.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust will
probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 November 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
33894 328/0124 97-585 PEM Power Off ZEPS
33897 328/0631 97-586 WINDII Power On Completed
33894 328/0124 97-587 OBC Clk Adj fr 41.2 to 20.6 msec/day
33898 328/0816 97-588 WINDII Power OFF
33903 328/1600 97-589 WINDII Power ON
33912 329/0645 97-590 OBC Clk Adj fr 20.6 to 41.2 msec/day
33921 329/2059 97-591 HRDI Power OFF, QB OFF
33922 329/2236 97-592 RTS New RTSUs 12 and 13 Loaded
33922 329/2236 97-593 PSU Comp HTRUs 19 and 20 Disabled
33924 330/0139 97-594 PEM Power On AXIS
33934 330/1715 97-595 HRDI HRDI to low power mode
33936 330/2026 97-596 HRDI HRDI htr delay to 25 amp-min
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 November 1997
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
OBSSOC2 Y-LO Battery 2 Current SOC 33866,880-2,86-87
PWTMPS1 Y-HI Mod Temp 1 33866,880,87
ACHTRI2 R-LO ACRIM Heater Current 2 33886-967
ACHTRV2 R-LO ACRIM Heater Voltage 2 33886-967
OBSBATVEON2 Y-LO Battery 2 voltage at EON 33896,99
OBSBATVEON3 Y-LO Battery 3 voltage at EON 33896,99
OBSVLBEON Y-LO Load Bus Volts at EON 33896,99
OBSBATVEON2 R-LO Battery 2 voltage at EON 33897-8,909-22
OBSBATVEON3 R-LO Battery 3 voltage at EON 33897-8,909-22
OBSVLBEON R-LO Load Bus Volts at EON 33897-8,909-22
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WIERRLIMIT WINDII Software Error Code 33896
WISRCBBS WINDII Source Broad Band 33904
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 November 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
33894 328/01:44 (24 Nov 97) from 41.2 to 20.6 msec/day
33912 329/06:51 (25 Nov 97) from 20.6 to 41.2 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 November 1997
22 Nov 1997 - GMT Day 326 Beta = 64.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.84/-4.94 +0.0/ +0.0 18.7 0.000 0.0
2 +11.2/ -22.4 +4.89/+3.23 +23.2/-16.0 25.9 1.020 20.9
3 +56.0/ -11.2 +3.70/+2.29 +22.4/-15.2 25.8 1.020 20.0
28 Nov 1997 - GMT Day 332 Beta = 77.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.94/-3.68 +0.0/ +0.0 18.7 0.000 0.0
2 +11.2/ -22.4 +2.98/+1.56 +16.8/-16.4 27.4 1.037 14.4
3 +22.4/ -22.4 +2.29/+0.94 +16.0/-16.0 27.2 1.040 13.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 November 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-119 FOT 33904 328/1739 14170 Generic Late Acquisition
97-119 FOT 33917 329/1523 14170 Generic Late Acquisition
97-122 FOT 33937 330/2239 N/A TAC Crashed
97-123 FOT 33945 331/1154 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period